# Memory Storyboard: Leveraging Temporal Segmentation for Streaming Self-Supervised Learning from Egocentric Videos

### v1

@@ 5.3 Other Training Factors
We experimented with a variation of Memory Storyboard as well as three baseline methods (SimCLR [11], Osiris [73], and Temporal Classification [40]) where the group normalization layers in the ResNet backbone are replaced with batch normalization [29] layers. The models are trained on SAYCam and evaluated on the downstream mini-ImageNet classification task with a SVM. The resulting accuracies are shown in Table 5. We observe that GroupNorm significantly outperform BatchNorm for all the models examined. This result is aligned the conclusion in [73] that BatchNorm is not compatible with unsupervised continual learning, and extends the conclusion to streaming SSL.
@@
Figure 5: Memory Storyboard model performance on SAYCam with different long-term memory sizes (5k, 10k, 50k, and 100k) and varying training batch compositions (12.5% – 75.0% from $M_{short}$) using SVM readout. Each colored line represents the performance of different training batch compositions when the model has seen the same amount of data from the stream. Each black line represent the performance of different training batch compositions when the model has taken the same number of gradient updates.
5.4 Optimal Batch Composition Under Different Memory Constraints
In Memory Storyboard, the training batch is composed of samples from both the long-term memory and the short-term memory (see Figure 2). However, the optimal composition ratio of the training batch, i.e. the optimal percentage of data in the training batch that comes from the short-term memory, is yet to be explored. Sampling more data from the short-term memory means we can digest more data within a fixed number of training steps, but there will be more distribution shift between different training batches. On the other hand, sampling more data from the long-term memory buffer may result in overfitting on the long-term memory data. In this section we experiment with different memory sizes and training composition, and demonstrate the optimal batch composition under different memory constraints.
We fix the size of the short-term memory $|M_{short}|$ to be $5K$, and vary the memory constraint for the long-term memory $|M_{long}|=5K,10K,50K,100K$. For each long-term memory size, we experiment with batch size from data stream $b=64,128,192,256,320,384$ (which corresponds to 12.5% though 75% of the training batch size). We sample $b$ images from the short-term memory and $512-b$ images from the long-term memory to compose a training batch. We evaluate the model with SVM readout on mini-ImageNet after the model has seen every 10% of the entire data stream and plot the results in Figure 5. We discuss the different observations for large memory size and small memory size respectively.
  • Large memory size. When the long-term memory size is large (Figures 5(c) and 5(d)), overfitting on the memory is unlikely and hence we can sample more data from the long-term memory and the performance still keeps increasing as the model sees more data. Hence, with the same amount of data seen by the model (colored curves), it is better to sample only a small batch from the short-term memory. However, when we control the number of model update steps to the same (black curves), neither focusing on the short-term memory or focusing on the long-term memory is preferable. In such case, the optimal batch size from the short term memory is at roughly 50% of the training batch.
  • Small memory size. When the long-term memory size is small (Figures 5(a) and 5(b)), the model is prone to overfitting on the memory. In our experiments, when at least 50% of the training batch (256 images) come from the short-term memory, the downstream task accuracy will start to decrease before we reach the end of the video stream. As a result, with the same number of model update steps (black curves), taking more images from the short-term memory gives better results. With the same amount of data seen by the model (colored curves), getting a higher percentage of long-term memory data has an advantage in the beginning when the memory size is not too small compared to the data seen by the model, but it is ultimately outperformed by models that focus more on short-term memory.
To summarize, the optimal training batch composition is dependent on the memory and compute constraint. A bigger batch from the long-term memory is preferred when we can afford a relatively large memory (in our experiments, 50K images from a 200-hour video stream) and we care about the model’s performance after seeing a fixed amount of data. A smaller batch from the long-term memory is preferred when we cannot afford to a large memory to prevent overfitting on the memory buffer data. When we can afford a large memory buffer and we care about the model’s performance after a fixed amount of computation for real-time learning, a balanced training batch composition is preferred.
6 Conclusion
The ability to continuously learn from large-scale uncurated streaming video data is crucial for applying self-supervised learning methods in real-world embodied agents. Existing works have limited exploration on this problem, have mainly focused on static datasets, and do not perform well in the streaming video setting. inspired by the event segmentation mechanism in human cognition, in this work, we propose Memory Storyboard, which leverages temporal segmentation to produce a two-tier memory hierarchy akin to the short-term and long-term memory of humans. Memory Storyboard combines a temporal contrastive objective and a standard self-supervised contrastive objective to facilitate representation learning from scratch through streaming video experiences. Memory Storyboard achieves state-of-the-art performance on downstream classification and object detection tasks when trained on real-world large egocentric video datasets. By studying the effects of subsampling rates, average segment length, normalization, and optimal batch composition under different compute and memory constraints, we also offer valuable insights on the design choices for streaming self-supervised learning.
Appendix A Experiment Details
Model Architecture
On top of the ResNet backbone, we use a two-layer MLP with 2048 hidden units, 128 output units, and ReLU activation function as the projector. In Memory Storyboard, we create two separate projectors for $\mathcal{L}_{TCL}$ and $\mathcal{L}_{CL}$.
Training
For all experiments in Tables 1 and 2, we used a total batch size of 512 (64 from $M_{short}$ and 448 from $M_{long}$ by default). The input resolution of the images to the model is 112. We apply a standard data augmentation pipeline for SSL methods following Zhuang et al., [75], which include random resized crop, random horizontal flip, random color jitter, random gray scale, random Gaussian filter, and color-normalization with ImageNet [13]. For the SimCLR [11], Osiris [73], and TC [40] experiments, we used the Adam [31] optimizer with a constant learning rate of 0.001, and a projector with 2 MLP layers of size 2048 and 128 respectively. For the SimSiam [12] experiments, we used the SGD optimizer with learning rate 0.05, momentum 0.9, and weight decay 1e-4, and a projector with 3 MLP layers of size 2048.
Evaluation
For mini-ImageNet and Labeled-S evaluations, the streaming SSL models are evaluated every 5% of the entire dataset. That is, we store 20 model checkpoints throughout the streaming training and evaluate them on mini-ImageNet and Labeled-S with SVM and kNN readout. The best result among these checkpoints are reported. Similar to Zhuang et al., [75], for SVM readout, we report the best performance among learning rate values {1e-7, 1e-6, 1e-5, 1e-4, 1e-3, 1e-2, 1e-1, 1, 1e1, 1e2}; for kNN readout, we report the best performance among $k$ values {1, 3, 5, 10, 20, 50, 100, 150, 200, 250}.
For ImageNet-1K evaluations, we evaluate the final model after streaming SSL training on the entire dataset. Following Purushwalkam et al., [43], we train a linear classifier on top of the normalized learned representations and report the classification accuracy. We used the LARS [66] optimizer with learning rate 3.0, momentum 0.9, and cosine learning rate schedule for 10 epochs. We used a batch size of 1024.
For OAK evaluations, we use Faster R-CNN [46], a popular two-stage object detector. We initialize the ResNet-50 [27] backbone with the backbone of the final checkpoint of the streaming SSL model, and fine-tune the entire model on OAK with IID training for 10 epochs, following the training configurations of [63].
Appendix B Additional Results
B.1 Separating Short-term Memory Batch and Long-term Memory Batch
@@
Inspired by the design of separating the loss on the new data and the replay data in Osiris [73], we investigate the optimal strategy of applying the temporal contrastive loss on the training batch. We consider applying the temporal contrastive loss only on data from short-term memory, only on data from long-term memory, separately on data from short-term and long-term memory and average the losses, and on the entire training batch (concatenated data from short-term and long-term memory). We report the results in Table 6. For experiments in the main paper, we apply the temporal contrastive loss only on data from long-term memory.
@@
The results here demonstrate that applying the temporal contrastive loss only on data from long-term memory or on the entire training batch achieve best performance. Applying the temporal contrastive loss only on data from short-term memory achieves inferior performance due to the limited number of temporal classes in the short-term buffer.
@@
Appendix C Optimal Batch Composition for SimCLR
@@
Figure 6: SimCLR model performance on SAYCam with different long-term memory sizes (5k, 10k, 50k, and 100k) and varying training batch compositions (12.5% – 75.0% from $M_{short}$) using SVM readout. Each colored line represents the performance of different training batch compositions when the model has seen the same amount of data from the stream. Each black line represent the performance of different training batch compositions when the model has taken the same number of gradient updates.
@@
We replicate the experiments in Figure 5 on SimCLR models with two-tier memory, and plot the results in Figure 6. We observe that the analysis and the conclusions of section 5.4 still holds: when we have a large memory, we either prefer balanced training batch (with fixed amount of computation) or a bigger batch from long-term memory (with fixed amount of data); when we can only afford a small memory, we prefer a smaller batch from long-term memory. We also want to note that, the SVM readout results starts to go down towards the end of the streaming training in SimCLR experiments more often than Memory Storyboard experiments, suggesting the better scalability of Memory Storyboard to larger-scale streaming training.
These results demonstrate that the analysis and observations in section 5.4 regarding the optimal batch composition for streaming SSL training under different memory and compute constraints is general, and applies to standard SSL methods in addition to Memory Storyboard.
Appendix D More Comprehensive Comparison between Memory Storyboard and SimCLR
With the experiment results in Figure 5 and Figure 6, we provide a more comprehensive comparision between Memory Storyboard and SimCLR performance under different memory constraints and batch compositions in Figure 7. We observe that Memory Storyboard outperforms SimCLR under the same amount of seen data, across a wide range of memory sizes and batch compositions. In particular, we note that Memory Storyboard significantly outperform SimCLR when we sample more data from $M_{short}$ (towards the right side of the $x$-axis). This results in the higher optimal performance when the memory size is small, where a larger batch from $M_{short}$ is needed to prevent overfitting on the long-term memory for better performance. We argue that, with temporal segmentation and the temporal contrastive loss, Memory Storyboard is able to provide better memory efficiency and also alleviate the temporal correlation issue suffered by SimCLR when we sample a large batch from the short-term memory.

### v3

@@ C.1 Performance of Different Normalization Layers
We experimented with a variation of Memory Storyboard as well as three baseline methods (SimCLR (Chen et al., 2020), Osiris (Zhang et al., 2024), and Temporal Classification (Orhan et al., 2020)) where the group normalization layers in the ResNet backbone are replaced with batch normalization (Ioffe & Szegedy, 2015) layers. The models are trained on SAYCam and evaluated on the downstream mini-ImageNet classification task with an SVM. The resulting accuracies are shown in Table 7. We observe that GroupNorm significantly outperforms BatchNorm for all the models examined. This result is aligned with the conclusion in (Zhang et al., 2024) that BatchNorm is less compatible with unsupervised continual learning, and extends the conclusion to streaming SSL.
@@
C.2 Separating Short-term Memory Batch and Long-term Memory Batch
@@
Inspired by the design of separating the loss on the new data and the replay data in Osiris (Zhang et al., 2024), we investigate the optimal strategy of applying the temporal contrastive loss on the training batch. We consider applying the temporal contrastive loss only on data from short-term memory, only on data from long-term memory, separately on data from short-term and long-term memory and average the losses, and on the entire training batch (concatenated data from short-term and long-term memory). We report the results in Table 8. For experiments in the main paper, we apply the temporal contrastive loss only on data from long-term memory.
@@
The results here demonstrate that applying the temporal contrastive loss only on data from long-term memory or on the entire training batch achieves the best performance. Applying the temporal contrastive loss only on data from short-term memory achieves inferior performance due to the limited number of temporal classes in the short-term buffer.
@@
C.3 Memory Storyboard in the IID Setting
To provide more context for the performance of Memory Storyboard in the Streaming Learning setting, we investigate the performance of Memory Storyboard in the IID setting. We take a SimCLR IID pre-trained model and use it to generate temporal segmentations on the entire dataset. Then we assign pseudo-labels to each frame according to the temporal segmentation results and train Memory Storyboard in the IID setting, taking the same number of gradient steps as the streaming setting. We observe that IID Memory Storyboard outperforms IID SimCLR and IID Simsiam, demonstrating the effectiveness of the temporal contrastive loss even in the IID setting.
Method SAYCam KrishnaCam
mini-INet INet iNat mini-INet INet iNat
SimCLR MemStoryboard (50K) 39.58 24.78 7.77 36.36 22.75 6.10
SimCLR MemStoryboard (50K) 41.32 26.37 9.85 35.20 22.75 6.71
IID SimCLR 44.04 30.44 8.69 36.90 23.77 5.60
IID SimSiam 29.02 20.92 4.91 28.58 22.28 4.16
IID SimCLR MemStoryboard 44.54 29.98 8.92 37.60 24.43 7.36
IID SimSiam MemStoryboard 42.30 31.02 10.07 36.54 26.60 7.06
Table 9: Performance of Memory Storyboard in the IID Setting.
C.4 Memory Storyboard with the Cross Entropy Loss
One alternative objective for the temporal contrastive loss in Memory Storyboard is the cross entropy (CE) loss. We investigate the performance of Memory Storyboard with the CE loss instead of the Supervised Contrastive (SupCon) loss. Results are summarized in Table 10. We observe that the CE loss outperforms the SupCon loss as the temporal contrastive loss when trained jointly with the SimCLR objective. We opted for the SupCon loss in the main text due to its flexibility. With the CE loss, we either need to know the number of temporal segments beforehand or gradually increase the size of the final classifier layer as the model sees more data, which is not ideal for streaming learning on a never-ending video stream. With the SupCon loss, we can learn from a very large number of temporal segments with the same model size.
Temp. Contrast Loss SAYCam 10K SAYCam 50K KrishnaCam 10K KrishnaCam 50K
mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat
SupCon 35.02 20.72 5.65 39.58 24.78 7.77 33.72 20.13 5.64 36.36 22.75 6.10
Cross Entropy 35.58 24.04 6.62 40.06 26.92 8.19 34.44 25.27 5.95 36.26 26.67 6.73
Table 10: Results on Memory Storyboard (using SimCLR as the base SSL method) with the cross entropy-loss instead of the supervised contrastive loss as the temporal contrastive loss.
C.5 Memory Storyboard with only the Temporal Contrastive Loss
To demonstrate the need for both the self-supervised loss and the temporal contrastive loss in our training objective, we experiment with using only the temporal contrastive (SupCon) loss and not the self-supervised loss (only using the self-supervised loss (SimCLR or SimSiam) and not the temporal contrastive loss has been experimented in the “two-tier” buffer baseline in Tables 1 and  2). Results are shown in Table 11. We observe that the performance of only using the SupCon loss is also inferior to the full memory storyboard method, demonstrating the necessity of joint training on both losses for best performance.
Method SAYCam 10K SAYCam 50K KrishnaCam 10K KrishnaCam 50K
mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat
SimCLR MemStoryboard 35.02 20.72 5.65 39.58 24.78 7.77 33.72 20.13 5.64 36.36 22.75 6.10
SimSiam MemStoryboard 36.72 22.99 6.66 41.32 26.37 9.85 33.78 21.38 6.51 35.20 22.75 6.71
Supcon Only 34.62 21.04 6.62 39.08 24.92 8.19 31.68 21.29 5.86 34.92 23.07 6.56
Table 11: Results on Memory Storyboard with only the temporal contrastive Loss.
C.6 Memory Storyboard with Multiple Gradient Steps per Batch
Using multiple gradient steps for each batch is a widely used technique in online continual learning Madaan et al. (2022). We investigate the performance of Memory Storyboard when we take multiple gradient steps on each batch. Results are shown in Table 12. We observe that using multiple gradient steps (2 or 4) produces a sizable improvement on the ImageNet readout evaluation on KrishnaCam but not on the other benchmarks. We also observed that the improvement of multiple gradient steps is a lot smaller on SAYCam (sometimes even harming the performance), presumably due to the fact that SAYCam is a much larger training dataset than KrishnaCam and streaming learning without multiple gradient steps is sufficient for the model to capture a wide range of visual concepts.
Method Grad SAYCam 10K SAYCam 50K KrishnaCam 10K KrishnaCam 50K
Steps mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat
SimCLR MemStoryboard 1 35.02 20.72 5.65 39.58 24.78 7.77 33.72 20.13 5.64 36.36 22.75 6.10
SimCLR MemStoryboard 2 34.56 23.47 5.66 38.44 26.49 7.67 33.00 23.86 5.45 35.30 26.36 6.38
SimCLR MemStoryboard 4 33.26 22.08 4.21 35.78 25.64 5.96 32.40 23.19 5.61 35.34 26.11 6.38
SimSiam MemStoryboard 1 36.72 22.99 6.66 41.32 26.37 9.85 33.78 21.38 6.51 35.20 22.75 6.71
SimSiam MemStoryboard 2 37.04 26.89 7.13 40.20 30.29 9.27 33.86 25.22 6.27 35.66 26.21 6.85
SimSiam MemStoryboard 4 35.02 22.88 6.73 36.30 25.21 7.36 33.44 24.34 6.44 34.98 25.73 6.62
Table 12: Results on Memory Storyboard with different number of gradient update steps per batch.
C.7 Memory Storyboard with Class-Balanced Buffer
Inspired by other methods with use smart memory storage policies (Yu et al., 2023; Purushwalkam et al., 2022), we investigate the performance of Memory Storyboard with a class-balanced memory. When we attempt to add a new data point to the long-term memory that is already full, we randomly remove one of the data points from the class with the most samples in the memory. Results are shown in Table 13. We observe that using the class-balanced memory produces mild improvements over the reservoir sampling baseline, though results on specific runs are mixed. We think that the memory storyboard method should work well with many different buffer sampling strategies, and advancements in buffer sampling strategies are orthogonal to the contribution of this work.
Base SSL Method Bal. Buffer SAYCam 10K SAYCam 50K KrishnaCam 10K KrishnaCam 50K
mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat mini-INet INet iNat
SimCLR 35.02 20.72 5.65 39.58 24.78 7.77 33.72 20.13 5.64 36.36 22.75 6.10
SimCLR 34.04 22.21 6.41 38.58 26.99 7.58 31.92 21.28 5.70 34.66 23.64 5.78
SimSiam 36.72 22.99 6.66 41.32 26.37 9.85 33.78 21.38 6.51 35.20 22.75 6.71
SimSiam 34.66 23.02 6.33 37.20 24.69 7.37 33.86 22.67 6.32 36.76 25.61 7.09
Table 13: Results on Memory Storyboard with class-balanced buffer.
C.8 Additional Qualitative Results
OAK Object Detection Results.
We visualize the object detection results produced by Memory Storyboard when fine-tuned on the OAK dataset (Wang et al., 2021) in Figure 5. We observe that the fine-tuned model can successfully detect objects in cluttered environments. The results show that the representations learned by Memory Storyboard can be effectively transferred to downstream tasks which requires more fine-grained features.
Label Merging Results.
We visualize the class labels produced by the label merging mechanism in Memory Storyboard in Figure 7. We observe that Memory Storyboard can successfully group semantically similar scenes together, which helps improve the representation learning performance.
Temporal Segmentation by Randomly Initialized Models.
We visualize the temporal segments produced by randomly initialized models in Figure 6. By comparing to Figure 3, we observe that randomly initialized models fail to capture intricate transitions between scenes and cannot create accurate temporal segments, while Memory Storyboard training enables the model to learn better image representations to capture more intricate scene transitions.
Figure 7: Visualization of label merging by Memory Storyboard on SAYCam. Each image represents a temporal segment; segments sharing the same color bar have been merged. Memory Storyboard successfully groups semantically similar scenes—e.g., segments marked with the light blue bar are all associated with the dining table.
Appendix D Optimal Batch Composition for SimCLR
@@
Figure 8: SimCLR model performance on SAYCam with different long-term memory sizes (5k, 10k, 50k, and 100k) and varying training batch compositions (12.5% – 75.0% from $M_{short}$) using SVM readout. Each colored line represents the performance of different training batch compositions when the model has seen the same amount of data from the stream. Each black line represents the performance of different training batch compositions when the model has taken the same number of gradient updates.
@@
We replicate the experiments in Figure 4 on SimCLR models with two-tier memory, and plot the results in Figure 8. We observe that the analysis and the conclusions of section 5.4 still hold: when we have a large memory, we either prefer balanced training batches (with a fixed amount of computation) or a bigger batch from long-term memory (with a fixed amount of data); when we can only afford a small memory, we prefer a smaller batch from long-term memory. We also want to note that the SVM readout results start to go down towards the end of the streaming training in SimCLR experiments more often than Memory Storyboard experiments, suggesting the better scalability of Memory Storyboard to larger-scale streaming training.
These results demonstrate that the analysis and observations in section 5.4 regarding the optimal batch composition for streaming SSL training under different memory and compute constraints are general, and apply to standard SSL methods in addition to Memory Storyboard.
Appendix E More Comprehensive Comparison between Memory Storyboard and SimCLR
With the experiment results in Figure 4 and Figure 8, we provide a more comprehensive comparison between Memory Storyboard and SimCLR performance under different memory constraints and batch compositions in Figure 9. We observe that Memory Storyboard outperforms SimCLR under the same amount of seen data, across a wide range of memory sizes and batch compositions. In particular, we note that Memory Storyboard significantly outperforms SimCLR when we sample more data from $M_{short}$ (towards the right side of the $x$-axis). This results in higher optimal performance when the memory size is small, where a larger batch from $M_{short}$ is needed to prevent overfitting on the long-term memory for better performance. We argue that, with temporal segmentation and the temporal contrastive loss, Memory Storyboard is able to provide better memory efficiency and also alleviate the temporal correlation issue suffered by SimCLR when we sample a large batch from the short-term memory.